6 people queue up to get on the bus. Is 2 specific person refuse to follow each other, how many ways are possible?
Found 2 solutions by stanbon, AlvinWong:
Answer by stanbon(75887)

6 people queue up to get on the bus. Is 2 specific person refuse to follow each other, how many ways are possible?
---
If the 2 were more flexible: 6! = 720 ways to line up
---
If the 2 always queued together: 2*5! = 240 ways to line up
-------------------------------
Problem condition # of lineups = 720-240 = 480 ways
=========================================================
Cheers,
Stan H.
=========

Answer by AlvinWong(1)

Use 6!-(5!x2!)=720-240=480